OpenStudy (shamil98):

Find the area of the region that lies inside the first curve and outside the second curve. r^2 = 8 cos 2theta , r = 2 so this is a leminscate

OpenStudy (shamil98):

4 = 8 cos 2theta 1/2 = cos 2theta theta = pi/6

OpenStudy (inkyvoyd):

dude r=2 is a circle

OpenStudy (shamil98):

oh right

OpenStudy (shamil98):

anyways i set it up but im off by a bit...

OpenStudy (shamil98):

\[\int\limits_{0}^{\pi/6} (8\cos2 \theta)^2 - (2)^2 d \theta\]

OpenStudy (shamil98):

@ganeshie8 what am i doing wrong here? i keep getting the wrong answer no matter how i set it up..

OpenStudy (shamil98):

\[1/2 \int\limits_{-\pi/6}^{\pi/6} 8\cos 2\theta - 4~ d \theta\]